OpenStudy (callisto):

Uncategorized question #1 Given that the decimal part of \[5+\sqrt{11}\] is A and the decimal part of \[5-\sqrt{11}\] is B. Let C=A+B, find the value of C.

OpenStudy (anonymous):

Decimal part is the fraction part right?

OpenStudy (experimentx):

1??

OpenStudy (callisto):

Fraction?! seems not... The answer I've written is 1.. I forgot how did I get the answer :|

OpenStudy (anonymous):

Then what is the decimal part?

OpenStudy (callisto):

The square root...

OpenStudy (anonymous):

I don't understand. What is the decimal part of 434.4545343434 ? 0.4545343434 isn't?

OpenStudy (callisto):

\[5+\sqrt{11} = 8.31662479....\] decimal part = 0.31662479....

OpenStudy (anonymous):

Yes that's the fractional part of the number.

OpenStudy (anonymous):

less than 1.

OpenStudy (callisto):

Okay... I don't know it can be expressed in a fraction :|

OpenStudy (callisto):

But how to solve it? Note: no calculator!

OpenStudy (anonymous):

No it can't be, actually programming knowledge terminology affected my answer.

OpenStudy (callisto):

Do you mean.... it is not solvable? Or...?

OpenStudy (callisto):

I think I can 'solve' it now... thanks.......

OpenStudy (anonymous):

It is, C is also a decimal number right?

OpenStudy (callisto):

C is 1....

OpenStudy (anonymous):

I am getting 1.28571 :/

OpenStudy (anonymous):

9/7 actually

OpenStudy (callisto):

._. \[5-\sqrt{11} = 1.68337521.....\] B= 0.68337521....

OpenStudy (callisto):

The solution is actually... quite simple :|

OpenStudy (anonymous):

:|

OpenStudy (callisto):

Should I post it?

OpenStudy (anonymous):

Yes mam, please :|

OpenStudy (callisto):

\[5+\sqrt{11}=5+3+A\]\[5−\sqrt{11}=1+B\]Add them.. \[10=(5+3+A)+(1+B)\]\[A+B=1\] This should be better :|

OpenStudy (experimentx):

Ah ... nice solution!! I had the intuition ... but couldn't think it would be so simple!!

OpenStudy (anonymous):

How \( 5−\sqrt{11}=1+B\) ? :|

OpenStudy (anonymous):

Got it :|

OpenStudy (callisto):

sqrt 11 = 3.xxxxxx 5-3.xxxxx = 1.yyyyy

OpenStudy (anonymous):

I am suck a fool :|

OpenStudy (callisto):

If fool = good, then you are :)

OpenStudy (anonymous):

NO, this was so easy :( I was thinking it too hard and in the lines of complicated things :|

OpenStudy (anonymous):

*such Good solution Calli~